what is the value of m if -5/8m-2=-12? can you show your work so i can process how it is done.
Answer by Theo(13342)

(-5/8) * m - 2 = -12

add 2 to both sides of the equation to get:

(-5/8) * m = 10

multiply both sides of the equation by 8 to get:

-5 * m = 80

divide both sides of the equation by -5 to get:

m = -16